3. Let X be a discrete random variable with the following PMF: 0.1 for x 0.2 for 0.2 for x=3 Pg(x)=〈 0.1 for x=4 0.25 for x=5 0.15 for x=6 otherwise a) (10 points) Find E[X] b) (10 points) Find Var(X) c) Let Y-* I. (15 points) Find E[Y] II. (15 points) Find Var(Y) X-HX 4. Consider a discrete random variable X with E [X]-4x and Var(X) = σ. Let Y a. (10 points) Find E[Y] b. (20 points) Find Var(Y)
Answer #1

3) from given pmf:

x P(x) xP(x) x2P(x)
1 0.1 0.1 0.1
2 0.2 0.4 0.8
3 0.2 0.6 1.8
4 0.1 0.4 1.6
5 0.25 1.25 6.25
6 0.15 0.9 5.4
total 1 3.65 15.95

a)E(X) =\sumxP(x) =3.65

b) E(X2) =\sumx2P(x) =15.95

hence Var(X) = E(X2)-(E(X))2 =15.95-3.652 =2.6275

c) from Y=2/X ;

below is distribution of Y:

y P(y) yP(y) y2P(y)
2.000 0.100 0.200 0.400
1.000 0.200 0.200 0.200
0.667 0.200 0.133 0.089
0.500 0.100 0.050 0.025
0.400 0.250 0.100 0.040
0.333 0.150 0.050 0.017
total 1 0.7333 0.7706

i) E(Y) =0.7333

ii) Var(Y) =0.7706-0.73332 =0.2328

4)

E(Y) =E((X- \mux)/\sigmax) =(1/\sigmax)*(E(X)-E( \mux)) =(1/\sigmax)*(\mux-\mux) =0

Var(Y) =Var((X- \mux)/\sigmax) =(1/\sigmax)*(Var(X)+Var( \mux)) =(1/\sigmax)*\sigmax =1